# Making Images Real Again: A Comprehensive Survey on Deep Image Composition

### v6

@@ I Introduction
Figure 3: Previous works perform multiple sub-tasks (e.g., object placement, image blending, image harmonization, shadow generation) sequentially or parallelly to achieve the goal of image composition.
TABLE I: The issues to be solved in image composition task and the corresponding deep learning methods to solve these issue. Note that some methods only focus on one issue while some methods attempt to solve multiple issues simultaneously. “Boundary” means refining the boundary between foreground and background. “Appearance” means adjusting the illumination of foreground. “Shadow” means generating shadow for the foreground. “Geometry” means seeking for reasonable location, size, and shape for the foreground considering geometric constraints. “Occlusion” means coping with the unreasonable occlusion. “Semantics” means finding suitable semantic context for the object.
boundary appearance shadow geometry occlusion semantics methods
@@
The appearance inconsistency is including but not limited to: 1) unnatural boundary between foreground and background; 2) incompatible illumination statistics between foreground and background; 3) missing or implausible shadow and reflection of foreground; 4) resolution, sharpness, and noise discrepancy between foreground and background [85]. For the first issue, the foreground is usually extracted using image segmentation [118] or matting [199, 36] algorithms. However, the foregrounds may not be precisely delineated, especially at the boundaries. When pasting the foreground with jagged boundaries on the background, there would be obvious color artifacts along the boundary. To solve this issue, image blending [191, 220] aims to address the unnatural boundary between foreground and background, so that the foreground could be seamlessly blended with the background. For the second issue, since the foreground and background may be captured in different conditions (e.g., weather, season, time of the day, camera setting), the obtained composite image could be inharmonious (e.g., foreground captured in the daytime and background captured at night). To solve this issue, image harmonization [170, 25, 21] aims to adjust the illumination statistics of foreground to make it more compatible with the background, so that the whole image looks more harmonious. For the third issue, when pasting the foreground on the background, the foreground may also affect the background with shadow or reflection. To solve this issue, shadow generation [101, 225, 145] or reflection generation [114] focus on generating plausible shadow or reflection for the foreground according to both foreground and background information. As far as we are concerned, there are only few works [114, 189, 165] on generating reflection for the inserted object probably due to the limited application scenarios, so we focus on shadow generation in this paper. For the fourth issue, the foreground and background may be from two images with different resolutions, blur degrees, and noise patterns. The resolution (resp., sharpness, noise) discrepancy between them could be mitigated by using super-resolution [186], deblurring [218], denoising [167] techniques.
@@
In the remainder of this paper, we will elaborate on each sub-task or combinatorial task. In particular, we will introduce object placement in Section II, image blending in Section III, image harmonization in Section IV, shadow generation in Section V, generative image composition in Section VI, foreground object search in Section VII. In each section, we will introduce the existing methods, available datasets, and common evaluation metrics. Finally, we will conclude the whole paper in Section VIII. The contributions of this paper can be summarized as follows,
@@ II-A Traditional Methods
Some object placement methods design explicit rules to find reasonable location and scale for the foreground object. For example, Remez et al. [137] proposed to move the foreground object of fixed scale along the same horizontal scanline on the background. They assume that the locations along the same horizontal scanline have similar depth, so that the true scale of foreground object can be well-preserved. Wang et al. [174] designed the instance-switching strategy to generate new images through switching different instances of the same class with similar shape and scale. To better refine the position where the object is pasted, Fang et al. [32] explored appearance consistency heatmap to guide the object placement, based on the intuition that an object could be moved to another location which has similar visual context to its original visual context. Specifically, one element in the appearance consistency heatmap measures the similarity between the visual context at this point and original visual context. Georgakis et al. [42] proposed to combine support surface detection and semantic segmentation to find proper location for placing the object. With the determined location, the size of the object is decided in the light of the depth at this location and the original scale of the object. Zhang et al. [226] proposed to model the probability distribution of bounding box information conditioned on background image and foreground category using Gaussian mixture model.
@@ II-B1 Category-specific Object Placement
Generative approaches: Tan et al. [157] proposed to predict the location and scale of inserted object by taking the background image and object layout as input. Besides, the bounding box prediction task is converted to a classification task by discretizing the locations and scales. Lee et al. [79] investigated on taking a background semantic map instead of a background image as input. Given a background semantic map, they designed a network consisting of two generative modules, in which the first module accounts for the bounding box of inserted object and the second module accounts for the mask shape of inserted object.
@@ II-B2 Instance-specific Object Placement
Generative approaches: Generative approaches [168, 219, 97, 90, 89, 204, 182] predict different types of spatial transformations (e.g., shifting and scaling, affine transformation, perspective transformation) for the foreground object, which is more flexible and powerful than category-specific object placement methods. For instance, Tripathi et al. [168] developed a model with generator, discriminator, and target network. Given a pair of background and foreground, the generator predicts the affine transformation for the foreground object to produce a composite image. The produced composite image is expected to fool the discriminator and fit the target network corresponding to a downstream task (e.g., object detection). Similarly, Zhan et al. [210] adopted spatial transformer network (STN) [63] to predict the warping parameters under an adversarial learning framework. To produce multiple reasonable placements, Zhang et al. [219] combined the foreground feature, background feature, and a random vector to predict the object placement. Moreover, they ensure the diversity of object placement by enforcing the pairwise distances between predicted placements to approach those between corresponding random vectors. To promote the diversity of generated placements, Zhou et al. [234] established the bijection between random vector and positive composite image. Moreover, they reformulated object placement as a graph completion task. In particular, background nodes have both content features and placements, while the inserted foreground node only has content feature, giving rise to an incomplete graph. Hence, they estimated the missing placement for the foreground node to complete the graph. Zhang et al. [224] proposed to make sequential decisions to produce a reasonable placement by using reinforcement learning. Qin et al. [134] employed a pre-trained large multi-modal model to generate a caption containing the placement information, and then predicted the placement bounding boxes.
@@
In terms of more advanced geometric transformation like view synthesis and pose transfer, some methods [2, 97, 214, 44] predicted perspective transformation to adjust the viewpoint and some methods [179, 203] predicted human pose in the scene context. Gou et al. [44] revealed that for perspective transformation, predicting the target locations of four source points is more effective than predicting the locations for more source points [214] or predicting the transformation parameters [97]. However, the view and pose synthesis ability is quite limited, especially for drastic viewpoint change (e.g., from front view to side view) and complicated human-object interaction (e.g., playing piano). To accomplish drastic viewpoint change and flexible pose transfer, generative composition methods attempted to re-generate the foreground object, which will be introduced in Section VI.
@@ II-C Datasets and Evaluation Metrics
In some previous works [168, 32], object placement is used as data augmentation strategy to facilitate the downstream tasks (e.g., object detection, instance segmentation). Therefore, they make use of existing object detection and instance segmentation datasets [98, 31, 24, 41]. In particular, the foregrounds are cropped out based on the annotated segmentation masks. After removing the foreground objects, the remaining incomplete background images are restored to complete background images by using image inpainting techniques [205, 102, 207]. In this manner, triplets of foregrounds, backgrounds, and ground-truth composite images can be obtained. Some other works focus on specific applications like 2D virtual try-on [97, 71, 90] (e.g., placing glasses/hats on human faces) or logo composition [89] (e.g., attaching logo to product image), so they need to collect foregrounds and backgrounds specifically for these applications. More recently, Liu et al. [103] released a large-scale object placement assessment dataset named OPA, which consists of 73,470 composite images and their binary rationality labels. OPA dataset is constructed by compositing the foregrounds and backgrounds from COCO dataset [98], followed by manually labelling the rationality of obtained composite images. A large number of annotated composite images could greatly facilitate the research on object placement. Qin et al. [134] established the OPAZ dataset following the format of OPA.
@@
To evaluate the quality of generated composite images, previous object placement works usually adopt the following three schemes: 1) Some works measure the similarity between real images and composite images. For example, Tan et al. [157] score the correlation between the distributions of predicted boxes and ground-truth boxes. Zhang et al. [219] calculate Frechet Inception Distance (FID) [54] between composite images and real images. However, they cannot evaluate each individual composite image. 2) Some works [168, 32] utilize the performance improvement of downstream tasks (e.g., object detection) to evaluate the quality of composite images, where the training sets of the downstream tasks are augmented with generated composite images. However, the evaluation cost is quite huge and the improvement in downstream tasks may not reliably reflect the quality of composite images, because it has been revealed in [43] that randomly generated unrealistic composite images could also boost the performance of downstream tasks. 3) Another common evaluation strategy is user study, where people are asked to score the rationality of object placement [79, 157]. User study complies with human perception and each composite image can be evaluated individually. However, due to the subjectivity of user study, the gauge in different papers may be dramatically different. There is no unified benchmark dataset and the results in different papers cannot be directly compared. 4) Finally, with the recently released OPA dataset [103], we could use the annotated composite images for evaluation. Nevertheless, the sparse annotations only cover a small proportion of locations and scales, which limits the universal evaluation of arbitrary composition results.
@@ III-B Deep Learning Methods
Inspired by traditional image blending methods [131, 7], some recent works [191, 220] explored incorporating the function of smoothing boundary into deep learning network. Among them, the works [191, 220, 216] not only enable smooth transition over the boundary, but also reduce the illumination discrepancy between foreground and background, in which the latter one is the goal of image harmonization in Section IV. In this section, we only introduce the way they enable smooth transition over the boundary. These two works [191, 220] are both inspired by [131]. Specifically, they add the gradient domain constraint to the objective function according to Poisson equation, which can produce a smooth blending boundary with gradient domain consistency. They both optimize over the input composite image to minimize the gradient domain loss. Differently, [191] has a close-form solution, while [220] converts the gradient domain loss to a differentiable loss function and uses gradient descent algorithm.
@@
More recently, mask-free image blending has emerged, which does not require initial masks. ControlCom [213] takes in a foreground image enclosing the foreground object and a background image with bounding box specifying the foreground placement, producing a composite image. The mask-free methods relieve the burden of initial mask prediction, which is not affected by the quality of initial masks. However, the shape of foreground object could be slightly altered and some details might be lost.
@@ III-C Datasets and Evaluation Metrics
The existing deep image blending works [191, 220, 216] adopt the following evaluation metrics: 1) calculating realism score using the pretrained model [235] which reflects the realism of a composite image; 2) conducting user study by asking engaged users to select the most realistic images; 3) Zhang et al. [216] deem the composite images obtained using ground-truth alpha matte as ground-truth composite image, and calculate Peak Signal-to-Noise Ratio (PSNR) between resultant image and ground-truth composite image.
@@ III-D Experiments
We evaluate different image blending methods conditioned on the matting results. First, we create composite images using the alpha matts predicted by the state-of-the-art trimap-based image matting methods [26, 107, 104]. Then, we hope that image blending methods can refine the obtained composite images. We sample 500 foreground images from recent image matting datasets [104, 84, 83]. For each foreground image, we randomly select two background images from BG20K [85]. The foreground images and background images form the test set.
@@ IV-A Rendering based Methods
Conventional image relighting [127, 148, 185, 221, 166] aims to adjust the appearance of an image or the object in an image as lit by novel illumination. With some adaptation, image relighting can also be used to adjust the foreground appearance according to the illumination of a new background [127, 221, 11, 59], which bears some resemblance to image harmonization. However, they usually achieve this goal by inferring explicit illumination condition, material properties, and 3D geometry, in which the supervision for these information is difficult and expensive to acquire. Besides, they generally have strong assumption for the light source, which may not generalize well to complicated real-world scenes.
@@ IV-B Non-rendering based Methods
Early traditional image harmonization methods [201, 155, 77, 151] is performing color transformation for the foreground to match the low-level color statistics between foreground and background. The difference between different methods mainly lies in the matching details. For example, Xue et al. [201] proposed to train a classifier to predict the zone (e.g., low, middle, high) of the histogram which can be best matched between foreground and background, and then adjust the foreground color to match the selected zone between foreground and background. [155] explored decomposing an image into a multi-resolution pyramid with multiple subbands, and performing histogram matching for each subband between foreground and background. Lalonde and Efros [77] proposed to represent foreground and background with color clusters, followed by matching foreground and background color clusters. Song et al. [151] proposed to calculate the color transformation (channel-wise scales) based on gray pixels of foreground and background, because normalized illumination color can be directly derived from the pixel values of gray pixels. In some works on sky replacement [238, 161, 169], they attempted to match the color statistics (e.g., mean, variance) between sky region and non-sky region. Broadly speaking, traditional color transfer methods [136, 195, 35, 132, 33, 1] can also be used for color matching between foreground and background to produce a harmonized image.
@@
Early deep learning based image harmonization methods target at making the harmonized images indistinguishable from real images. For instance, Zhu et al. [235] explored predicting the realism of an image using a CNN classifier. With such realism predictor, they learn the color transformation for the foreground to achieve high realism score, and also enforce the color variation in different channels to be close. Similar to [235], the works [211, 13] used adversarial learning to make the harmonized images indistinguishable from real images. Bhattad and Forsyth [5] drew inspiration from Retinex theory [72] that an image can be decomposed into albedo (reflectance) and shading (illumination). On the premise of this assumption, an image harmonization model is trained so that the harmonized result should have consistent albedo and consistent background shading with input composite image.
@@
With the emergence of image harmonization datasets consisting of paired training data (see Section IV-E), abundant image harmonization methods [170, 21, 23, 106, 111, 158, 16, 206, 184, 130] using paired supervision have been developed. Tsai et al. [170] proposed the first end-to-end CNN network for image harmonization and leveraged auxiliary semantic segmentation branch to enhance the basic image harmonization network. Another work Sofiiuk et al. [150] also utilized high-level semantic features, which are inserted into the encoder to provide auxiliary information. Cun and Pun [25] designed an additional Spatial-Separated Attention Module to deal with foreground and background feature maps separately. Hao et al. [52] employed self-attention [180] mechanism to propagate relevant information from background to foreground. By treating different capture conditions as different domains, Cong et al. [21] proposed a domain verification discriminator to pull close the foreground domain and background domain. Similarly, Cong et al. [22] formulated image harmonization as background-guided domain translation task, in which the domain code of background is directly used to guide the harmonization process. One byproduct of [22] is predicting the inharmony level of an image by comparing the domain codes of foreground and background, so that we can selectively harmonize those apparently inharmonious composite images. Inspired by [22], Valanarasu et al. [171] proposed to extract style code from part of background.
@@
Recently, some diffusion-based image harmonization models [92, 17, 233, 232, 138] have applied conditional diffusion model to image harmonization task. Zhou et al. [232] proposed to modify VAE decoder to alleviate the image distortion issue of diffusion model. Ren et al. [138] proposed to inject background illumination information into diffusion model.

### v7

@@ I Introduction
Figure 3: Previous works perform multiple sub-tasks (e.g., object placement, image blending, image harmonization, shadow/reflection generation) sequentially or parallelly to achieve the goal of image composition.
TABLE I: The issues to be solved in image composition task and the corresponding deep learning methods to solve these issue. Note that some methods only focus on one issue while some methods attempt to solve multiple issues simultaneously. “Boundary” means refining the boundary between foreground and background. “Appearance” means adjusting the illumination of foreground. “Shadow” means generating shadow for the foreground. “Reflection” means generating reflection for the foreground. “Geometry” means seeking for reasonable location, size, and shape for the foreground considering geometric constraints. “Occlusion” means coping with the unreasonable occlusion. “Semantics” means finding suitable semantic context for the object.
boundary appearance shadow reflection geometry occlusion semantics methods
@@
+ + + + + + + [95]
The appearance inconsistency is including but not limited to: 1) unnatural boundary between foreground and background; 2) incompatible illumination statistics between foreground and background; 3) missing or implausible shadow and reflection of foreground; 4) resolution, sharpness, and noise discrepancy between foreground and background [94]. For the first issue, the foreground is usually extracted using image segmentation [129] or matting [218, 42] algorithms. However, the foregrounds may not be precisely delineated, especially at the boundaries. When pasting the foreground with jagged boundaries on the background, there would be obvious color artifacts along the boundary. To solve this issue, image blending [210, 241] aims to address the unnatural boundary between foreground and background, so that the foreground could be seamlessly blended with the background. For the second issue, since the foreground and background may be captured in different conditions (e.g., weather, season, time of the day, camera setting), the obtained composite image could be inharmonious (e.g., foreground captured in the daytime and background captured at night). To solve this issue, image harmonization [185, 29, 25] aims to adjust the illumination statistics of foreground to make it more compatible with the background, so that the whole image looks more harmonious. For the third issue, when pasting the foreground on the background, the foreground may also affect the background with shadow or reflection. To solve this issue, shadow generation [111, 248, 158] or reflection generation [125, 208, 180] focus on generating plausible shadow or reflection for the foreground according to both foreground and background information. For the fourth issue, the foreground and background may be from two images with different resolutions, blur degrees, and noise patterns. The resolution (resp., sharpness, noise) discrepancy between them could be mitigated by using super-resolution [204], deblurring [239], denoising [182] techniques.
@@
In the remainder of this paper, we will elaborate on each sub-task or combinatorial task. In particular, we will introduce object placement in Section II, image blending in Section III, image harmonization in Section IV, shadow (resp., reflection) generation in Section V (resp., VI), generative image composition in Section VII, foreground object search in Section VIII. In each section, we will introduce the existing methods, available datasets, and common evaluation metrics. Finally, we will conclude the whole paper in Section IX. The contributions of this paper can be summarized as follows,
@@ II-A Traditional Methods
Some object placement methods designed explicit rules to find reasonable location and scale for the foreground object. For example, Remez et al. [149] proposed to move the foreground object of fixed scale along the same horizontal scanline on the background. They assumed that the locations along the same horizontal scanline have similar depth, so that the true scale of foreground object can be well-preserved. Wang et al. [189] designed the instance-switching strategy to generate new images through switching different instances of the same class with similar shape and scale. To better refine the position where the object is pasted, Fang et al. [38] explored appearance consistency heatmap to guide the object placement, based on the intuition that an object could be moved to another location which has similar visual context to its original visual context. Specifically, one element in the appearance consistency heatmap measures the similarity between the visual context at this point and original visual context. Georgakis et al. [48] proposed to combine support surface detection and semantic segmentation to find proper location for placing the object. With the determined location, the size of the object is decided in the light of the depth at this location and the original scale of the object. Zhang et al. [249] proposed to model the probability distribution of bounding box information conditioned on background image and foreground category using Gaussian mixture model.
@@ II-B1 Category-specific Object Placement
Generative approaches: Tan et al. [172] proposed to predict the location and scale of inserted object by taking the background image and object layout as input. Besides, the bounding box prediction task is converted to a classification task by discretizing the locations and scales. Lee et al. [87] investigated on taking a background semantic map instead of a background image as input. Given a background semantic map, they designed a network consisting of two generative modules, in which the first module accounts for the bounding box of inserted object and the second module accounts for the mask shape of inserted object. Parihar et al. [139] focused on “person” category and leveraged the prior knowledge of text-to-image generation model. Specifically, the foreground mask and foreground image are jointly optimized to fit the background and text prompt, after which the foreground mask can be used to indicate the person placement.
@@ II-B2 Instance-specific Object Placement
Generative approaches: Generative approaches [183, 240, 107, 100, 99, 224, 200] predict different types of spatial transformations (e.g., shifting and scaling, affine transformation, perspective transformation) for the foreground object, which is more flexible and powerful than category-specific object placement methods. For instance, Tripathi et al. [183] developed a model with generator, discriminator, and target network. Given a pair of background and foreground, the generator predicts the affine transformation for the foreground object to produce a composite image. The produced composite image is expected to fool the discriminator and fit the target network corresponding to a downstream task (e.g., object detection). Similarly, Zhan et al. [231] adopted spatial transformer network (STN) [70] to predict the warping parameters under an adversarial learning framework. To produce multiple reasonable placements, Zhang et al. [240] combined the foreground feature, background feature, and a random vector to predict the object placement. Moreover, they ensured the diversity of object placement by enforcing the pairwise distances between predicted placements to approach those between corresponding random vectors. To promote the diversity of generated placements, Zhou et al. [259] established the bijection between random vector and positive composite image. Moreover, they reformulated object placement as a graph completion task. In particular, background nodes have both content features and placements, while the inserted foreground node only has content feature, giving rise to an incomplete graph. Hence, they estimated the missing placement for the foreground node to complete the graph. Zhang et al. [246] proposed to make sequential decisions to produce a reasonable placement by using reinforcement learning. Qin et al. [146] employed a pre-trained large multi-modal model to generate a caption containing the placement information, and then predicted the placement bounding boxes. Zhou et al. [255] first predicted candidate bounding boxes and then associated each object query with matched bounding box.
@@
In terms of more advanced geometric transformation like view synthesis and pose transfer, some methods [3, 107, 235, 50] predicted perspective transformation to adjust the viewpoint and some methods [197, 223] predicted human pose in the scene context. Gou et al. [50] revealed that for perspective transformation, predicting the target locations of four source points is more effective than predicting the locations for more source points [235] or predicting the transformation parameters [107]. However, the view and pose synthesis ability is quite limited, especially for drastic viewpoint change (e.g., from front view to side view) and complicated human-object interaction (e.g., playing piano). To accomplish drastic viewpoint change and flexible pose transfer, generative composition methods attempted to re-generate the foreground object, which will be introduced in Section VII.
@@ II-C Datasets and Evaluation Metrics
In some early works [183, 38], object placement is used as data augmentation strategy to facilitate the downstream tasks (e.g., object detection, instance segmentation). Therefore, they make use of existing object detection and instance segmentation datasets [108, 37, 28, 47]. In particular, the foregrounds are cropped out based on the annotated segmentation masks. After removing the foreground objects, the remaining incomplete background images are restored to complete background images by using image inpainting techniques [225, 112, 227]. In this manner, triplets of foregrounds, backgrounds, and ground-truth composite images can be obtained. Some other works focus on specific applications like 2D virtual try-on [107, 78, 100] (e.g., placing glasses/hats on human faces) or logo composition [99] (e.g., attaching logo to product image), so they need to collect foregrounds and backgrounds specifically for these applications. More recently, Liu et al. [113] released a large-scale object placement assessment dataset named OPA, which consists of 73,470 composite images and their binary rationality labels. OPA dataset is constructed by compositing the foregrounds and backgrounds from COCO dataset [108], followed by manually labelling the rationality of obtained composite images. A large number of annotated composite images could greatly facilitate the research on object placement. Qin et al. [146] established the OPAZ dataset following the format of OPA.
@@
To evaluate the quality of generated composite images, previous object placement works usually adopt the following schemes: 1) Some works measure the similarity between real images and composite images. For example, Tan et al. [172] score the correlation between the distributions of predicted boxes and ground-truth boxes. Zhang et al. [240] calculate Frechet Inception Distance (FID) [60] between composite images and real images. However, they cannot evaluate each individual composite image. 2) Some works [183, 38] utilize the performance improvement of downstream tasks (e.g., object detection) to evaluate the quality of composite images, where the training sets of the downstream tasks are augmented with generated composite images. However, the evaluation cost is quite huge and the improvement in downstream tasks may not reliably reflect the quality of composite images, because it has been revealed in [49] that randomly generated unrealistic composite images could also boost the performance of downstream tasks. 3) Another common evaluation strategy is user study, where people are asked to score the rationality of object placement [87, 172]. User study complies with human perception and each composite image can be evaluated individually. OPA dataset [113] has annotated composite images and its test set could be used for evaluation. Nevertheless, the sparse annotations only cover a small proportion of locations and scales, which limits the universal evaluation of arbitrary composition results. 4) To support the evaluation of arbitrary composition results, we can train a binary classifier based on annotated positive and negative composite images, to predict the rationality score of an arbitrary composite image.
@@ II-D Experiments
For practical usage, we recommend fast discriminative approaches [132, 262] which are more stable, effective, and flexible.
@@ III-B Deep Learning Methods
Inspired by traditional image blending methods [143, 8], some recent works [210, 241] explored incorporating the function of smoothing boundary into deep learning network. Among them, the works [210, 241, 237] not only enabled smooth transition over the boundary, but also reduced the illumination discrepancy between foreground and background, in which the latter one is the goal of image harmonization in Section IV. In this section, we only introduce the way they enable smooth transition over the boundary. These two works [210, 241] are both inspired by [143]. Specifically, they added the gradient domain constraint to the objective function according to Poisson equation, which can produce a smooth blending boundary with gradient domain consistency. They both optimized over the input composite image to minimize the gradient domain loss. Differently, [210] had a close-form solution, while [241] converted the gradient domain loss to a differentiable loss function and uses gradient descent algorithm.
@@
More recently, mask-free image blending has emerged, which does not require initial masks. ControlCom [234] received a foreground image enclosing the foreground object and a background image with bounding box specifying the foreground placement, producing a composite image. The mask-free methods relieve the burden of initial mask prediction, which is not affected by the quality of initial masks. However, the shape of foreground object could be slightly altered and some details might be lost.
@@ III-C Datasets and Evaluation Metrics
The existing deep image blending works [210, 241, 237] adopt the following evaluation metrics: 1) Zhang et al. [237] deem the composite images obtained using ground-truth alpha matte as ground-truth composite image, and calculate Peak Signal-to-Noise Ratio (PSNR) between resultant image and ground-truth composite image. 2) conducting user study by asking engaged users to select the most realistic images; 3) calculating realism score using the pretrained model [260] which reflects the realism of a composite image.
@@ III-D Experiments
We evaluate different image blending methods conditioned on the matting results. First, we create composite images using the alpha mattes predicted by the state-of-the-art trimap-based image matting methods [30, 118, 114]. Then, we hope that image blending methods can refine the obtained composite images. We sample 500 foreground images from recent image matting datasets [114, 93, 92]. For each foreground image, we randomly select two background images from BG20K [94]. The foreground images and background images form the test set.
@@
For practical usage, traditional image blending methods are recommended under specific circumstances. For example, when the foreground mask is accurate and there is no high demand for sharp boundary, alpha blending [145] could be used. When the foreground has the same target color as surrounding background, Poisson blending [143] could be used. In other complex cases, learnable image blending methods [237, 234] are recommended, but they may struggle in dealing with transparent foreground and delicate edges. In the cases of transparent foreground (e.g., glass) and delicate edges (e.g., floating hair strands), it is suggested to improve the performance of image matting methods.
@@ IV-A Rendering based Methods
Conventional image relighting [138, 162, 203, 242, 181] aims to adjust the appearance of an image or the object in an image as lit by novel illumination. With some adaptation, image relighting can also be used to adjust the foreground appearance according to the illumination of a new background [138, 242, 12, 65], which bears some resemblance to image harmonization. However, they usually achieve this goal by inferring explicit illumination condition, material properties, or 3D geometry, in which the supervision for these information is difficult and expensive to acquire. Besides, they generally have strong assumption for the light source, which may not generalize well to complicated real-world scenes.
@@ IV-B Non-rendering based Methods
Early traditional image harmonization methods [221, 170, 85, 165] performed color transformation for the foreground to match the low-level color statistics between foreground and background. The difference between different methods mainly lies in the matching details. For example, Xue et al. [221] proposed to train a classifier to predict the zone (e.g., low, middle, high) of the histogram which can be best matched between foreground and background, and then adjust the foreground color to match the selected zone between foreground and background. [170] explored decomposing an image into a multi-resolution pyramid with multiple subbands, and performing histogram matching for each subband between foreground and background. Lalonde and Efros [85] proposed to represent foreground and background with color clusters, followed by matching foreground and background color clusters. Song et al. [165] proposed to calculate the color transformation (channel-wise scales) based on gray pixels of foreground and background, because normalized illumination color can be directly derived from the pixel values of gray pixels. In some works on sky replacement [263, 176, 184], they attempted to match the color statistics (e.g., mean, variance) between sky region and non-sky region. Broadly speaking, traditional color transfer methods [148, 214, 41, 144, 39, 2] can also be used for color matching between foreground and background to produce a harmonized image.
@@
Early deep learning based image harmonization methods target at making the harmonized images indistinguishable from real images. For instance, Zhu et al. [260] explored predicting the realism of an image using a CNN classifier. With such realism predictor, they learned the color transformation for the foreground to achieve high realism score, and also enforced the color variation in different channels to be close. Similar to [260], the works [232, 15] used adversarial learning to make the harmonized images indistinguishable from real images. Bhattad and Forsyth [6] drew inspiration from Retinex theory [80] that an image can be decomposed into albedo (reflectance) and shading (illumination). On the premise of this assumption, an image harmonization model is trained so that the harmonized result should have consistent albedo and consistent background shading with input composite image.
@@
With the emergence of image harmonization datasets consisting of paired training data (see Section IV-E), abundant image harmonization methods [185, 25, 27, 116, 122, 173, 18, 226, 202, 142, 190] using paired supervision have been developed. Tsai et al. [185] proposed the first end-to-end CNN network for image harmonization and leveraged auxiliary semantic segmentation branch to enhance the basic image harmonization network. Another work Sofiiuk et al. [164] also utilized high-level semantic features, which are inserted into the encoder to provide auxiliary information. Cun and Pun [29] designed an additional Spatial-Separated Attention Module to deal with foreground and background feature maps separately. Hao et al. [58] employed self-attention [198] mechanism to propagate relevant information from background to foreground. By treating different capture conditions as different domains, Cong et al. [25] proposed a domain verification discriminator to pull close the foreground domain and background domain. Similarly, Cong et al. [26] formulated image harmonization as background-guided domain translation task, in which the domain code of background is directly used to guide the harmonization process. One byproduct of [26] is predicting the inharmony level of an image by comparing the domain codes of foreground and background, so that we can selectively harmonize those apparently inharmonious composite images. Inspired by [26], Valanarasu et al. [186] proposed to extract style code from part of background.
@@
Recently, some diffusion-based image harmonization models [91, 19, 258, 257, 150, 244, 117] have applied conditional diffusion model to image harmonization task. Zhou et al. [257] proposed to modify VAE decoder to alleviate the image distortion issue of diffusion model. Ren et al. [150] proposed to inject background illumination information into diffusion model.